OpenStudy (kj4uts):

A breeding group of beavers is introduced into a protected area. After t years the number of beavers in the area is modeled by the function N(t)= 54/0.35+0.68^t 1. How many beavers were initially introduced? 2. Estimate the number of beavers after 5 years. 3. Determine the change in the beaver population between t = 5 and t = 10. (Note: All answers are whole numbers.)

OpenStudy (kj4uts):

The answers that I have so far are (I am not sure if I did them right): 1. N(t)= 54/0.35+0.68^0 = 40 2. N(t)= 54/0.35+0.68^5 = 109 3. I am not sure?

OpenStudy (anonymous):

For initial result put t=0

OpenStudy (michele_laino):

question #1 you have to evaluate N(0)

OpenStudy (anonymous):

yup and for #2 put t=5

OpenStudy (kj4uts):

1. 40 2. 109 3. ?

OpenStudy (anonymous):

as change is n(5) - n(10) put t= 5 and t=10 and subtract

OpenStudy (michele_laino):

yes! for third part we have to compute this: \[{N\left( 5 \right) - N\left( 0 \right) = ...?}\]

OpenStudy (kj4uts):

I am not sure exactly how to compute that what do I plug in?

OpenStudy (michele_laino):

is your formula like this: \[N\left( t \right) = \frac{{54}}{{0.35 + 0.68t}}\]

OpenStudy (anonymous):

from 10 to 5 n(10) = 54/(.35+(.68^(10)))

OpenStudy (kj4uts):

@Michele_Laino that's how my formula is.

OpenStudy (michele_laino):

sorry ! Is it like below: \[\Large N\left( t \right) = \frac{{54}}{{0.35 + {{0.68}^t}}}\]

OpenStudy (kj4uts):

yes

OpenStudy (anonymous):

yup ^ this power operator.

OpenStudy (michele_laino):

then if t=0, we have: \[\Large N\left( 0 \right) = \frac{{54}}{{0.35 + 1}} = ...?\]

OpenStudy (kj4uts):

40

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

if t=5, then we have: \[\Large N\left( 5 \right) = \frac{{54}}{{0.35 + {{0.68}^5}}} \cong 109\]

OpenStudy (kj4uts):

so do I just go 109-40=69?

OpenStudy (anonymous):

nope i dont think so

OpenStudy (michele_laino):

for t=10, we have: \[\Large N\left( {10} \right) = \frac{{54}}{{0.35 + {{0.68}^{10}}}} \cong 145.5\]

OpenStudy (anonymous):

3. Determine the change in the beaver population between t = 5 and t = 10. here t is 10 and 5 implies 145-109 =?

OpenStudy (michele_laino):

so you have to do this: \[\Large N\left( {10} \right) - N\left( 5 \right) = 145.5 - 109 = ...?\]

OpenStudy (kj4uts):

145.5-109=36.5

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yup in whole no.

OpenStudy (michele_laino):

yes! @AdityaOS is right!

OpenStudy (kj4uts):

So the answers are: 1. 40 2. 109 3. 36.5 but isn't 36.5 not in whole number

OpenStudy (michele_laino):

you can round off 36.5 to 37

OpenStudy (kj4uts):

oh ok thank you @Michele_Laino and @AdityaOS

OpenStudy (michele_laino):

:)